Suppose we have a set of blood pressures with a mean of 70 Diastolic, and a sample standard deviation of 10 points. Assuming a normal distribution of Diastolic blood pressures, between what two values should 95% of all Diastolic blood pressures lie?
Answer by stanbon(75887)

Suppose we have a set of blood pressures with a mean of 70 Diastolic, and a sample standard deviation of 10 points. Assuming a normal distribution of Diastolic blood pressures, between what two values should 95% of all Diastolic blood pressures lie?
-----
Find the z-value with a left tail of 0.025
invNorm(0.025) = -1.96
Note: Correspondingly z = 1.96 has a right tail of 0.025
---
Find the corresponding blood prssure values using x = z*s + u
----
lower value: x = -1.96*10 + 70 = 50.4
upper value: x = +1.96*10 + 70 = 89.6
=========================================
Cheers,
Stan H.